# Degradation-Aware All-in-One Image Restoration via Latent Prior Encoding

### v2

@@ 4.7 Ablation Study
We conducted an ablation study to evaluate DAIR’s key components: Latent Priors Encoding (LPE) (”which”), Degradation Map (DM) (”where”), Latent fusion (LF) (”what”), and Restoration Block (3WD). As shown in Table 6, the inclusion of the proposed component progressively enhances performance, with DAIR achieving the highest average PSNR of 26.90 and SSIM of 0.8565. Notably, removing Latent Priors results in significantly worse performance (-4.23 PSNR / -0.0876 SSIM), highlighting its critical role in capturing detailed degradation descriptions, subsequently guiding the model in effectively segregating different types of degradations.
Table 6: Ablation study on key components. Tick () indicates the component is used, cross () indicates not used. Best results in bold red.
Method LP LF DM 3WD Denoise Desnowing Derain Lowlight Average
PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑
Base Model 23.42/0.7237 20.76/0.7696 22.70/0.7698 13.99/0.5111 20.22/0.6936
Base + LP 24.50/0.7417 20.75/0.7783 22.46/0.7658 14.23/0.5240 20.49/0.7025
Base + LP + LF 25.83/0.7799 21.21/0.8014 22.23/0.7545 15.04/0.5474 21.08/0.7208
Base + LP + LF + DM 26.80/0.8158 21.97/0.8442 22.79/0.7614 15.52/0.5540 21.77/0.7439
DAIR w/o Latent Prior 26.07/0.8855 26.83/0.8306 22.76/0.7741 15.00/0.5853 22.67/0.7689
DAIR 29.11/0.8971 31.59/0.9581 30.21/0.9070 16.68/0.6638 26.90/0.8565
Improvement +5.69/+0.1734 +10.83/+0.1885 +7.51/+0.1372 +2.69/+0.1527 +6.68/+0.1629
@@ A.2 VAE Design
To effectively integrate the VAE into our framework, we design a hybrid U-Net encoder–decoder with multi-head self-attention Tian et al. (2025a) at the bottleneck. In particular, both the hybrid VAE and the reconstruction network incorporate the same spatial feature dimensions, enabling seamless interaction between latent representations and restoration features. The encoder progressively downsamples RGB images through four stages with channel dimensions $[40,80,160,320]$ using 3×3 convolutions (stride=2 for downsampling, padding=1). Each stage contains a Residual Attention Block He et al. (2016) combining spatial convolutions with channel attention Hu et al. (2018) (reduction ratio 8):
@@
The KL divergence Kullback & Leibler (1951) term regularizes the latent distribution:
@@
To encourage discriminative latent representations, we incorporate weakly Supervised Contrastive Learning Guo et al. (2023) on the spatially-pooled encoder:
@@ A.4 Training Details
Appendix B Additional Experimental Results
B.1 Quantitative Comparison
B.1.1 Low-light
@@
We utilize the LSD dataset Sharif et al. (2025), collected in uncontrolled low-light settings, offering diverse indoor and outdoor scenes under varying conditions. Table 7 compares image restoration methods across extreme lowlight (under 50 Lux) and lowlight scenarios (50-200 lux), divided into indoor and outdoor subsets, using PSNR and SSIM metrics. Our method, DAIR, consistently outperforms others, with the best results highlighted in bold red and the second-best underlined. DAIR achieves significant gains, improving up to +2.64 PSNR and +0.0737 SSIM in extreme lowlight, showcasing its robustness and superior generalization in challenging conditions.
@@
Uformer Wang et al. (2022) 11.98/0.5278 13.24/0.4628 10.39/0.5535 12.79/0.4145
Restormer Zamir et al. (2022) 14.99/0.6343 16.35/0.5828 16.07/0.7121 13.90/0.5429
AIRNet Li et al. (2022) 8.51/0.3678 9.80/0.3248 8.13/0.4585 9.05/0.3107
PromptIR Vaishnav et al. (2023) 14.74/0.5902 16.10/0.5038 14.59/0.6404 13.48/0.4154
DiffUIR Zheng et al. (2024) 9.77/0.5185 11.86/0.3713 9.77/0.5185 11.29/0.3193
ADAIR Cui et al. (2024) 12.92/0.5456 14.46/0.4783 10.93/0.5576 12.72/0.4079
DFPIR Tian et al. (2025a) 13.32/0.5643 15.68/0.4815 15.63/0.6599 14.74/0.3871
@@
B.1.2 Denoising
Our experiment assesses the generalization of image denoising models under variable noise conditions, addressing limitations of prior works that train on fixed noise levels, leading to overfitting. We retrained baseline methods and our DAIR model using random Gaussian noise on DIV2K Agustsson & Timofte (2017) and evaluated on unseen datasets, BSD100 Martin et al. (2001) and Urban100 Kim et al. (2016). Table 8 illustrate performance comparison across these datasets and noise levels ($\sigma$ = 15, 25, 50). Notably, baseline models showed inconsistent performance; Restormer struggled at higher noise, DiffUIR Zheng et al. (2024) excelled on BSD100 at $\sigma$=50 but failed elsewhere, on six-task common AIR settings. Notably, MP-based methods, such as PromptIR Vaishnav et al. (2023), underperformed in the absence of precise noise-level prompts. This inconsistent performance, particularly among MP-based methods, underscores their inability to effectively segregate degradations, even on homogeneous tasks. In contrast, DAIR consistently achieved higher scores, with improvements of up to +2.2 dB PSNR at high noise levels, highlighting its superior robustness and ability to effectively understand degradations, thereby validating the effectiveness of its latent prior encoding.
@@ B.1.2 Denoising
Uformer Wang et al. (2022) 31.71/0.9265 29.16/0.8753 25.46/0.7351 33.28/0.9219 30.61/0.8635 33.28/0.9219 31.52/0.9387 29.13/0.8944 25.15/0.7664
Restormer Zamir et al. (2022) 27.38/0.8567 23.44/0.6825 16.62/0.3646 28.32/0.8385 23.73/0.6373 31.59/0.8937 28.11/0.8873 23.67/0.7304 16.52/0.4315
AIRNet Li et al. (2022) 32.91/0.9377 30.52/0.9095 27.07/0.8307 34.20/0.9333 31.83/0.8986 27.03/0.7006 33.19/0.9506 31.00/0.9324 27.46/0.8665
PromptIR Vaishnav et al. (2023) 28.84/0.8464 25.90/0.7306 21.45/0.5164 30.04/0.8295 27.03/0.7007 30.61/0.8636 29.23/0.8755 26.05/0.7734 21.23/0.5685
DiffUIR Zheng et al. (2024) 19.92/0.8065 15.95/0.6774 14.34/0.5435 21.78/0.7909 17.41/0.6917 31.99/0.9122 18.28/0.7967 14.86/0.6701 13.67/0.5496
ADAIR Cui et al. (2024) 32.51/0.9412 30.10/0.9075 26.80/0.8174 34.06/0.9354 31.59/0.8937 14.07/0.4531 32.74/0.9514 30.33/0.9233 26.72/0.8420
DFPIR Tian et al. (2025a) 22.13/0.8095 21.57/0.7585 20.43/0.6234 21.77/0.8164 21.47/0.7535 21.48/0.7533 22.36/0.8182 21.80/0.7759 21.48/0.7533
@@
B.2 More Visual Results
Fig. 7 presents visual results for denoising, deraining, and dehazing tasks, comparing our method with ADAIR Cui et al. (2024), DFPIR Tian et al. (2025a), degraded inputs, and reference images. These examples demonstrate that our approach consistently achieves clearer and more accurate restorations, closely matching the reference images and outperforming baselines across all degradation types.
@@ B.2 More Visual Results
Fig. 8 presents more visual results for compound degradations (e.g., low-light + haze + rain, haze + rain), comparing our method with ADAIR Cui et al. (2024), DFPIR Tian et al. (2025a), degraded inputs, and reference images. Our approach consistently delivers clearer and more accurate restorations, outperforming baselines in challenging compound degradation settings.
@@ B.3 Inference Benchmarking
Uformer Wang et al. (2022) 31.07 32.19 0.53 58.08 17.22 1.13 203.83 4.91 1.29
PromptIR Vaishnav et al. (2023) 65.62 15.24 0.25 180.31 5.55 0.36 707.17 1.41 0.37
ADAIR Cui et al. (2024) 71.01 14.08 0.23 191.50 5.22 0.34 732.94 1.36 0.36
DFPIR (with CLIP) Tian et al. (2025a) 70.08 14.27 0.23 184.33 5.43 0.36 714.20 1.40 0.37
DFPIR (dummy text) Tian et al. (2025a) 67.88 14.73 0.24 184.12 5.43 0.36 714.38 1.40 0.37
@@
Table 9 presents the inference speed of image restoration models, evaluated on lower-mid hardware (RTX 3060) using full-precision weights (FP32). Our DAIR model consistently achieves the fastest processing times, highest FPS, and superior throughput across various image sizes, outperforming all baseline AIR models. Notably, DAIR matches or surpasses the single-task model like Uformer (43.86G) Wang et al. (2022), which is significantly lighter compared to the existing AIR methods. These results highlight DAIR’s efficiency, demonstrating its ability to deliver SOTA restoration quality while maintaining faster speed and practicality for real-world applications, even on modest hardware setups.
B.4 Downstream Vision Tasks (Object Detection)
A primary objective of AIR is enhancing downstream vision task performance Cui et al. (2024); Jiang et al. (2025), yet existing AIR methods lack benchmarking on such tasks due to insufficient annotated datasets. To address this critical gap, we created a comprehensive evaluation framework by annotating 3,000 test images across six common degradations (lowlight Sharif et al. (2025), deraining Fu et al. (2017), dehazing Li et al. (2018), desnowing Li et al. (2020), denoising Agustsson & Timofte (2017), and deblurring Nah et al. (2017)). Our annotation process leveraged the paired nature of these datasets by first annotating ground truth reference images and transferring these annotations to their corresponding corrupted counterparts, ensuring consistency and reliability. We then enhanced these degraded images using several baseline AIR restoration methods and systematically evaluated their performance using the YOLOv12-Large Tian et al. (2025b) object detection model. This novel benchmarking approach provides the first quantitative assessment of how different restoration techniques impact object detection performance, offering valuable insights for developing AIR methods optimized for real-world computer vision applications.

### v3

@@ 4.6 Ablation Study
Table 9: Ablation study on key components. Tick () indicates the component is used, cross () indicates not used.
Method LP LF DM 3WD PSNR↑/SSIM↑
Base Model 20.22/0.6936
Base + LP 20.49/0.7025
Base + LP + LF 21.08/0.7208
Base + LP + LF + DM 21.77/0.7439
DAIR w/o Latent Prior 22.67/0.7689
DAIR 26.90/0.8565
Improvement +6.68/+0.1629
We conducted an ablation study combining four challenging tasks (denoising, desnowing, deraining, LLIE) to evaluate the key components of DAIR: Latent Priors Encoding (LPE) (”which”), Degradation Map (DM) (”where”), Latent Fusion (LF) (”what”), and Restoration Block (3WD). As shown in Table 9, the inclusion of the proposed component progressively enhances performance, with DAIR achieving the highest average PSNR of 26.90 and SSIM of 0.8565. Notably, removing Latent Priors results in significantly worse performance (-4.23 PSNR / -0.0876 SSIM), highlighting its critical role in capturing detailed degradation descriptions, subsequently guiding the model in effectively segregating different types of degradations. Please see Appendix B for module-wise ablation and details.
4.7 Real-world Implications
Table 10: Object Detec-
tion on restored images.
Method Average Precision
AP50:95 AP50 AP75
PromptIR 30.9 33.3 31.9
DFPIR 27.9 30.0 29.3
ADAIR 28.9 31.2 30.2
DAIR (Ours) 34.9 37.3 36.4
Improvement +4.0 +4.0 +4.5
Table 11: Real-world unseen restoration.
Method NIQE↓/MUSIQ↑/BRISQUE↓
PromptIR 9.65/43.81/40.03
DFPIR 8.01/45.02/25.52
ADAIR 6.85/46.10/28.45
DAIR (Ours) 5.25/50.19/25.97
Improvement -1.60/+4.09/+0.45
Tables 11 and 11 evaluate DAIR’s restoration quality through downstream tasks and real-world unseen restoration. For OD (Table 11), YOLOv12 tested on degraded MS-COCO images shows DAIR achieves 34.9 AP50:95, 37.3 AP50, and 36.4 AP75, outperforming PromptIR by +4.0, +4.0, and +4.5 points, demonstrating preserved semantic content. For perceptual quality (Table 11), no-reference metrics across real-world unseen datasets show DAIR achieves best scores: 5.25 NIQE (-1.60), 50.19 MUSIQ (+4.09), and 25.97 BRISQUE (+0.45), confirming superior perceptual quality and strong generalization to unseen out-of-distribution scenarios. See Appendix D for details.
@@ A.2 VAE Design
To effectively integrate the VAE into our framework, we design a hybrid U-Net encoder–decoder with multi-head self-attention (tian2025degradation) at the bottleneck. In particular, both the hybrid VAE and the reconstruction network incorporate the same spatial feature dimensions, enabling seamless interaction between latent representations and restoration features. The encoder progressively downsamples RGB images through four stages with channel dimensions $[40,80,160,320]$ using 3×3 convolutions (stride=2 for downsampling, padding=1). Each stage contains a Residual Attention Block (he2016resnet) combining spatial convolutions with channel attention (hu2018squeeze) (reduction ratio 8):
@@
The KL divergence (kullback1951information) term regularizes the latent distribution:
@@
To encourage discriminative latent representations, we incorporate weakly Supervised Contrastive Learning (guo2023letting) on the spatially-pooled encoder:
@@
Capture degradation with VAE.
Fig. 8 illustrates the proposed VAE’s progressive convergence and degradation-aware representation learning (Step = 0 → 100k). Incorporating SupCon loss as weak supervision improves latent separability of corruption types while preserving reconstruction fidelity, enabling the model to disentangle mixed degradations and generalize beyond the training distribution. Notably, it separates the unseen “Low + Rain” (guo2024onerestore) from the related seen “Low + Haze + Rain,” and characterizes out-of-domain cases such as underwater images, evidencing robust clustering and stable training dynamics.
Figure 8: Learning degradation with VAE. (a) VAE without SupCon. (b) VAE at Step = 0. (c) VAE at Step = 100k. (d) VAE (fully trained).
@@ A.4 Training Details
Appendix B Analysis and Ablation of DAIR
We conduct an ablation and module-wise analysis on four common tasks: denoising, desnowing, deraining, and LLIE. This section provides a detail the performance and contributions of our proposed module across these tasks.
B.1 Latent Prior Encoding (”which”)
We analyze the “which” component and demonstrate that LC separation makes feature selection both explicit and effective. By routing structure-dominant cues through luminance and color-specific cues through chrominance, the encoder preserves complementary statistics that can be modulated by latent priors per branch. This design validates our reasoning by enabling degradation-aware, stage-wise selection of “which” features to leverage, resulting in substantial gains over luminance-only, chrominance-only, and RGB based encoders, as shown in Table 12.
Table 12: Ablation study of which module.
Method PSNR↑/SSIM↑/LPIPS↓
Luma only 23.58/0.7996/0.1946
Chroma only 23.41/0.7927/0.2014
RGB 23.44/0.8002/0.1892
Proposed (Both) 26.90/0.8565/0.1316
Improvement +3.32/+0.0569/-0.0630
B.2 Degradation Map (”where”)
We evaluate the impact of the degradation map (”where”) separately. Table 13 demonstrates that the proposed degradation map with the FFT module significantly improves restoration performance. The model without the degradation map achieved a PSNR of 22.80, SSIM of 0.7769, and LPIPS of 0.2116, whereas our approach with the degradation map improved these metrics to a PSNR of 26.90, SSIM of 0.8565, and LPIPS of 0.1316, resulting in enhancements of +4.10 in PSNR, +0.0796 in SSIM, and a reduction of -0.0800 in LPIPS. It also ensures that the degradation map captures frequency-domain degradation characteristics, leveraging frequency cues.
Table 13: Ablation study of degradation map (”where”) module.
Method PSNR↑/SSIM↑/LPIPS↓
w/o FFT 22.80/0.7769/0.2116
Proposed 26.90/0.8565/0.1316
Improvement +4.10/+0.0796/-0.0800
B.3 Latent Fusion (”what”)
We systematically evaluate the impact of different fusion strategies on restoration performance. Table 14 illustrates the results of latent fusion across various settings. Our full $\mu$-fusion (Factor=1) achieves a PSNR of 26.90 dB, outperforming the second-best method by +3.92 dB. Notably, removing degradation-aware components (beta: 22.97 dB, MU-fusion: 22.74 dB) or employing fixed fusion factors (Factor=0: 17.88 dB) results in significant performance drops. This underscores that the learned $\mu$-fusion weights effectively capture degradation characteristics and facilitate semantically meaningful decisions.
Table 14: Ablation study of fusion strategies. Best in bold red, second underlined.
Method Beta Alpha Factor PSNR↑/SSIM↑/LPIPS↓
w/o MU 22.74/0.7791/0.2135
w/o beta 1 22.97/0.7824/0.2147
alpha + beta 0 17.88/0.7272/0.2488
Fusion 0.1 0.1 16.28/0.7350/0.2345
Fusion 0.5 0.5 22.98/0.7837/0.2143
Proposed 1 26.90/0.8565/0.1316
Improvement +3.92/+0.0728/-0.0827
B.4 Detail Ablation
We provide a module-wise detailed ablation study of the proposed DAIR in Table 15. This table illustrates the contributions of key components to the overall restoration performance. Each row represents a different configuration, with tick marks () indicating the inclusion of specific components and crosses () indicating their exclusion. The best results are highlighted in bold red. Notably, the full DAIR model demonstrates superior performance across all metrics, validating the effectiveness of the integrated components.
Table 15: Ablation study on key components.
Method LP LF DM 3WD Denoise Desnowing Derain Lowlight
PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑
Base Model 23.42/0.724 20.76/0.770 22.70/0.770 13.99/0.511
Base + LP 24.50/0.742 20.75/0.778 22.46/0.766 14.23/0.524
Base + LP + LF 25.83/0.780 21.21/0.801 22.23/0.755 15.04/0.547
Base + LP + LF + DM 26.80/0.816 21.97/0.844 22.79/0.761 15.52/0.554
DAIR w/o Latent Prior 26.07/0.886 26.83/0.831 22.76/0.774 15.00/0.585
DAIR 29.11/0.897 31.59/0.958 30.21/0.907 16.68/0.664
Appendix C Additional Comparison with SOTA Methods
C.1 Comparison on Three-task Settings
Our primary motivation is to automate the prompting process and reframe AIR for practical, real-world applications. In real-world scenarios, degraded scenes are often highly complex and frequently involve overlapping degradations. To demonstrate the effectiveness of our method in addressing both homogeneous and heterogeneous degradations, we evaluate its practicality in a challenging six-task heterogeneous setting comprising complex datasets (LSD for real-world lowlight (sharif2025illuminatingdarknesslearningenhance), random denoising instead of fixed denoising, heavy rain streak, etc.). We refer to these tasks as a common task in the manuscript. Furthermore, we incorporate a compound 5D task, where degradations are similar and overlapping (e.g., Haze + Snow, Haze + Lowlight + Snow). Thus, we can illustrate the limitations of existing work and clearly articulate our motivation. However, in recent times, many methods, including PromptIR, ADAIR, DFPIR, Perceive-IR, etc., have leveraged three task settings to illustrate restoration performance. Therefore, to establish a clear positioning of our proposed method relative to existing models, we benchmarked it under similar settings. Table 16 compares our DAIR with existing methods. Our proposed method achieves notable improvements of +1.13 dB in dehazing and +0.28 dB on average over the best baselines. These results underscore the effectiveness of DAIR in enhancing restoration quality while being significantly lightweight.
Table 16: Comparison of image restoration methods across different tasks and noise levels. Best results in bold red, second best underlined, and increment over best baseline highlighted in blue.
Method Dehazing Deraining Denoising (CBSD68) Average
SOTS Rain100L $\sigma=15$ $\sigma=25$ $\sigma=50$
PromptIR 30.58/0.974 36.37/0.972 33.98/0.933 31.31/0.888 28.06/0.799 32.06/0.913
Restormer 30.43/0.975 36.55/0.975 33.84/0.931 31.18/0.885 27.90/0.790 31.98/0.911
DFPIR 31.87/0.980 38.65/0.982 34.14/0.935 31.47/0.893 28.25/0.806 32.88/0.919
ADAIR 31.06/0.980 38.64/0.983 34.12/0.935 31.45/0.892 28.19/0.802 32.69/0.918
DAIR(Ours) 32.80/0.981 38.86/0.983 34.31/0.934 31.60/0.893 28.22/0.807 33.16/0.920
Improvement +0.93/+0.001 +0.22/+0.000 +0.17/-0.001 +0.13/+0.000 -0.03/+0.001 +0.28/+0.001
C.2 Common Task Details
C.2.1 Low-light
@@
We utilize the LSD dataset (sharif2025illuminatingdarknesslearningenhance), collected in uncontrolled low-light settings, offering diverse indoor and outdoor scenes under varying conditions. Table 17 compares image restoration methods across extreme lowlight (under 50 Lux) and lowlight scenarios (50-200 lux), divided into indoor and outdoor subsets, using PSNR and SSIM metrics. Our method, DAIR, consistently outperforms others, with the best results highlighted in bold red and the second-best underlined. DAIR achieves significant gains, improving up to +2.64 PSNR and +0.0737 SSIM in extreme lowlight, showcasing its robustness and superior generalization in challenging conditions.
@@
Uformer 11.98/0.5278 13.24/0.4628 10.39/0.5535 12.79/0.4145
Restormer 14.99/0.6343 16.35/0.5828 16.07/0.7121 13.90/0.5429
AIRNet 8.51/0.3678 9.80/0.3248 8.13/0.4585 9.05/0.3107
PromptIR 14.74/0.5902 16.10/0.5038 14.59/0.6404 13.48/0.4154
DiffUIR 9.77/0.5185 11.86/0.3713 9.77/0.5185 11.29/0.3193
ADAIR 12.92/0.5456 14.46/0.4783 10.93/0.5576 12.72/0.4079
DFPIR 13.32/0.5643 15.68/0.4815 15.63/0.6599 14.74/0.3871
@@
C.2.2 Denoising
Our experiment assesses the generalization of image denoising models under variable noise conditions, addressing limitations of prior works that train on fixed noise levels, leading to overfitting. We retrained baseline methods and our DAIR model using random Gaussian noise on DIV2K (agustsson2017div2k) and evaluated on unseen datasets, BSD100 (martin2001bsd) and Urban100 (kim2016vdsr). Table 18 illustrate performance comparison across these datasets and noise levels ($\sigma$ = 15, 25, 50). Notably, baseline models showed inconsistent performance; Restormer struggled at higher noise, DiffUIR (zheng2024selective) excelled on BSD100 at $\sigma$=50 but failed elsewhere, on six-task common AIR settings. Notably, MP-based methods, such as DFPIR (tian2025degradation), underperformed in the absence of precise noise-level prompts. This inconsistent performance, particularly among MP-based methods, underscores their inability to effectively segregate degradations, even on homogeneous tasks. In contrast, DAIR consistently achieved higher scores, with improvements of up to +2.2 dB PSNR at high noise levels, highlighting its superior robustness and ability to effectively understand degradations, thereby validating the effectiveness of its latent prior encoding.
@@ C.2.2 Denoising
Uformer 31.71/0.9265 29.16/0.8753 25.46/0.7351 33.28/0.9219 30.61/0.8635 33.28/0.9219 31.52/0.9387 29.13/0.8944 25.15/0.7664
Restormer 27.38/0.8567 23.44/0.6825 16.62/0.3646 28.32/0.8385 23.73/0.6373 31.59/0.8937 28.11/0.8873 23.67/0.7304 16.52/0.4315
AIRNet 32.91/0.9377 30.52/0.9095 27.07/0.8307 34.20/0.9333 31.83/0.8986 27.03/0.7006 33.19/0.9506 31.00/0.9324 27.46/0.8665
PromptIR 28.84/0.8464 25.90/0.7306 21.45/0.5164 30.04/0.8295 27.03/0.7007 30.61/0.8636 29.23/0.8755 26.05/0.7734 21.23/0.5685
DiffUIR 19.92/0.8065 15.95/0.6774 14.34/0.5435 21.78/0.7909 17.41/0.6917 31.99/0.9122 18.28/0.7967 14.86/0.6701 13.67/0.5496
ADAIR 32.51/0.9412 30.10/0.9075 26.80/0.8174 34.06/0.9354 31.59/0.8937 14.07/0.4531 32.74/0.9514 30.33/0.9233 26.72/0.8420
DFPIR 22.13/0.8095 21.57/0.7585 20.43/0.6234 21.77/0.8164 21.47/0.7535 21.48/0.7533 22.36/0.8182 21.80/0.7759 21.48/0.7533
@@
C.3 More Visual Results
Fig. 9 presents visual results for denoising, deraining, and dehazing tasks, comparing our method with ADAIR (cui2024adair), DFPIR (tian2025degradation), degraded inputs, and reference images. These examples demonstrate that our approach consistently achieves clearer and more accurate restorations, closely matching the reference images and outperforming baselines across all degradation types.
@@ C.3 More Visual Results
Fig. 10 presents more visual results for compound degradations (e.g., low-light + haze + rain, haze + rain), comparing our method with ADAIR (cui2024adair), DFPIR (tian2025degradation), degraded inputs, and reference images. Our approach consistently delivers clearer and more accurate restorations, outperforming baselines in challenging compound degradation settings.
@@ C.4 Inference Benchmarking
Uformer 31.07 32.19 0.53 58.08 17.22 1.13 203.83 4.91 1.29
PromptIR 65.62 15.24 0.25 180.31 5.55 0.36 707.17 1.41 0.37
ADAIR 71.01 14.08 0.23 191.50 5.22 0.34 732.94 1.36 0.36
DFPIR (with CLIP) 70.08 14.27 0.23 184.33 5.43 0.36 714.20 1.40 0.37
DFPIR (dummy text) 67.88 14.73 0.24 184.12 5.43 0.36 714.38 1.40 0.37
@@
Table 19 presents the inference speed of image restoration models, evaluated on lower-mid hardware (RTX 3060) using full-precision weights (FP32). Our DAIR model consistently achieves the fastest processing times, highest FPS, and superior throughput across various image sizes, outperforming all baseline AIR models. Notably, DAIR matches or surpasses the single-task model like Uformer (43.86G) (wang2022uformer), which is significantly lighter compared to the existing AIR methods. These results highlight DAIR’s efficiency, demonstrating its ability to deliver SOTA restoration quality while maintaining faster speed and practicality for real-world applications, even on modest hardware setups.
Appendix D More Results from Real-world
D.1 Downstream Vision Tasks (OD)
A primary objective of AIR is enhancing downstream vision task performance (cui2024adair; jiang2025survey), yet existing AIR methods lack benchmarking on such tasks due to insufficient annotated datasets. To address this critical gap, we created a comprehensive evaluation framework by annotating 3,000 test images across six common degradations (lowlight (sharif2025illuminatingdarknesslearningenhance), deraining (fu2017deep), dehazing (li2018benchmarking), desnowing (li2020desnownet), denoising (agustsson2017div2k), and deblurring (nah2017deep)). Our annotation process leveraged the paired nature of these datasets by first annotating ground truth reference images and transferring these annotations to their corresponding corrupted counterparts, ensuring consistency and reliability. We then enhanced these degraded images using several baseline AIR restoration methods and systematically evaluated their performance using the YOLOv12-Large (tian2025yolov12) OD model. This novel benchmarking approach provides the first quantitative assessment of how different restoration techniques impact OD performance, offering valuable insights for developing AIR methods optimized for real-world computer vision applications.